OpenStudy (anonymous):

In major league baseball, the pitcher's mound is 60 feet from the batter. If a pitcher throws a 91 mph fastball, how much time elapses from when the ball leaves the pitcher's hand until the ball reaches the batter?

OpenStudy (anonymous):

you get it or no

OpenStudy (anonymous):

hello

OpenStudy (campbell_st):

convert 91 mph to feet per second. the use the formula Time = Distance x speed

OpenStudy (anonymous):

I got .66

OpenStudy (anonymous):

got 5460

OpenStudy (campbell_st):

well its \[feet/\sec = \frac{91 \times 5280}{60 \times 60}\] when you get the feet per second you can use Time = speed/distance sorry about changing the equation, I just has a rethink on it... but Time = distance/speed will give the amount of time you have to react

OpenStudy (anonymous):

thank you for the example

OpenStudy (anonymous):

still don't get it

OpenStudy (anonymous):

were did you get the other 60 from

OpenStudy (campbell_st):

ok... so the baseball travels 133.466666... feet per second how long does it take to travel 60 ft time = 60/133.466666...

OpenStudy (campbell_st):

hope it makes sense

OpenStudy (anonymous):

iam good at this but it just cant get the right answer

OpenStudy (campbell_st):

well what did you get..?

OpenStudy (campbell_st):

are you happen the speed is 133.46666... feet per second..?

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

but were did you get the other 60

OpenStudy (campbell_st):

ok... so you are being asked how long does it take to travel 60 ft

OpenStudy (campbell_st):

so time = distance / speed

OpenStudy (anonymous):

ok thanks if I have any more question ill ask you

OpenStudy (campbell_st):

it should be between 0.4 and 0.5 seconds... at a guess

OpenStudy (anonymous):

it say it wrong